A ball falls from height H.There are n collisions with the Earth if after n collisions the ball Rises to a height y then coefficient of restitution is given by?

Grade:12th pass

1 Answers

Anhiti Rao
6 years ago
Coefficient of restitution when an objects hits/bounces off a stationary object/platform ise^2=y/Hafter n collisions it will become (e^n)^2=y/HTherefore coff. Restitution = (e)^2n=y/H or H×((e)^2n)=y